Please help me solve this word problem:
Tony is 10 years older than Sam. In five years Tony will be twice as old as Sam. How old are they now?
I need to solve the word problem using two equations. Here's what I get.
x=Sam's age Y=Tony's age
x+10=y
5+Y=2X

Answer by jorel1380(3719)

Tony=Sam+10
T=S+10
T+5=2(S+5)
S+10+5=2S+10
S+15=2S+10
5=S
Sam is 5. Tony is 15.